# Combinatorics and Hodge theory of degenerations of abelian varieties: A survey of the Mumford construction

### v2

@@ 3. The Mumford construction
In Section 2.5, we have shown how to extend ${\mathcal{A}}_{g}\to{\mathcal{A}}_{g}^{\,{\mathbb{B}}}$ (or $\widetilde{{\mathcal{A}}}_{g}\to\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$) toroidally, so that the period map $(\Delta^{*})^{k}\to{\mathcal{A}}_{g}$ of any degeneration $f\colon X\to\Delta^{k}$ with monodromy cone ${\mathbb{B}}_{f}={\mathbb{B}}$ admits an extension of the period map over the punctures $\Delta^{k}\to{\mathcal{A}}_{g}^{{\mathfrak{F}}}$. We will now describe how to extend the universal family of PPAVs over ${\mathcal{A}}_{g}$ so that we may pull back this extension, to produce particularly nice birational models of degenerations. It is useful to have examples in mind; many are provided in Section 3.5. All degenerations we consider in this section are maximal, in the sense of Definition 2.13.
@@ 3.1. Mumford construction, fan version
(We must take $|u|$ small enough that $B\log|u|-\log|a_{ij}|>0$.) Then this construction may be performed relatively, choosing $a$ in some subtorus of ${\rm Sym}_{g\times g}({\mathbb{C}}^{*})$ forming coset representatives for the natural action of $u\in{\mathbb{C}}^{*}$. We get an analytic degeneration
$f^{\rm univ}\colon X^{\rm univ}({\rm Cone}({\mathcal{T}}))\to\Delta^{\rm univ}$
where $\Delta^{\rm univ}\to{\rm Sym}_{g\times g}({\mathbb{C}}^{*})/{\mathbb{C}}^{*}$ is a holomorphic disk bundle over ${\rm Sym}_{g\times g}({\mathbb{C}}^{*})/{\mathbb{C}}^{*}$, such that $f^{\rm univ}$ restricts to a family of PPAVs over the punctured disc bundle $(\Delta^{\ast})^{\rm{univ}}\subset\Delta^{\rm{univ}}$.
In terms of the toroidal extensions of Section 2.5, $\Delta^{\rm univ}$ maps to the toroidal extension ${\mathcal{A}}_{g}^{{\mathfrak{F}}}$ whose fan is (the orbit of) a single ray ${\mathfrak{F}}=\operatorname{GL}_{g}({\mathbb{Z}})\cdot{\mathbb{R}}_{\geq 0}B$. So every $1$-parameter degeneration of PPAVs with monodromy $B$ admits an extension pulled back from $f^{\rm univ}$ along an arc transverse to the boundary divisor $\{0\}\times{\rm Sym}_{g\times g}({\mathbb{C}}^{*})/{\mathbb{C}}^{*}=\Delta^{\rm univ}\setminus(\Delta^{*})^{\rm univ}$. $\clubsuit$
@@
When $\mathcal{T}$} has integral vertices, the total space $X({\rm Cone}({\mathcal{T}}))$ or $X^{\rm univ}({\rm Cone}({\mathcal{T}}))$ of the Mumford degeneration is smooth if and only if $\mathcal{T}$} is a complete triangulation, i.e. $\mathcal{T}$} is a tiling by standard lattice simplices. This condition ensures that ${\rm Cone}({\mathcal{T}})$ is a regular fan—its cones are all standard affine cones. Then $X({\rm Cone}({\mathcal{T}}))\to\Delta$ is a semistable, $K$-trivial degeneration (with smooth total space), see Proposition 3.13 below. These are sometimes called Kulikov models, in analogy to K3 degenerations.
@@
By passing to an intermediate cover ${\mathcal{H}}_{g}\to\widetilde{{\mathcal{A}}}_{g}\to{\mathcal{A}}_{g}$ as in Remark 2.28, we may assume that the Levi quotient $\Gamma\subset\operatorname{GL}_{g}({\mathbb{Z}})$ of the parabolic stabilizer at some Baily–Borel 0-cusp of $\widetilde{{\mathcal{A}}}_{g}$ acts freely on ${\mathcal{P}}_{g}$. Then, when $B$ is primitive, the boundary divisor of $\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{R}}_{\geq 0}B}$ is isomorphic to the discriminant divisor of $f^{\rm univ}$, rather than a further finite quotient of it. In this case, the base of $f^{\rm univ}$ glues onto $\widetilde{{\mathcal{A}}}_{g}$ and thus, we may view $f^{\rm univ}$ as an extension of the universal family over $\widetilde{{\mathcal{A}}}_{g}$ to a family over $\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{R}}_{\geq 0}B}$. A priori, this extension exists only in the category of analytic spaces. But it is always an algebraic space (Prop. 6.12), and for certain choices of tiling $\mathcal{T}$}, we may ensure it is relatively projective, see Section 3.3.
When $B$ is not primitive, the base $\Delta^{\rm univ}$ of $f^{\rm univ}$ rather maps to $\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{R}}_{\geq 0}B}$ by a map ramified over the toroidal boundary divisor, to order the imprimitivity of $B$.
@@
satisfy $\widetilde{\Phi}(\tau)=\widetilde{\Phi}_{\rm nilp}(\tau)=\tau B,$ where $u=e^{2\pi i\tau}$ is the coordinate on $\Delta^{*}$. More generally, the period map on the restriction of $f^{\rm univ}$ to the universal cover of $(\Delta^{*})^{\rm univ}|_{\{a\}}$ for $a\in{\rm Sym}_{g\times g}({\mathbb{C}}^{*})$, is given by $\widetilde{\Phi}(\tau)=\widetilde{\Phi}_{\rm nilp}(\tau)=\tfrac{1}{2\pi i}(\log a_{ij})+\tau B$.
@@
The fan $\mathcal{S}$} must then be $\bf M$}-periodic (with respect to the action (12)) and the projection to ${\mathbb{R}}^{k}$ must induce a morphism of fans to $({\mathbb{R}}_{\geq 0})^{k}$, such that ${\rm Supp}({\mathcal{S}})$ contains ${\mathbb{R}}^{g}\times(1,\dots,1)$. Furthermore, we should require that the morphism ${\mathcal{S}}\to({\mathbb{R}}_{\geq 0})^{k}$ is flat, that is, the image of any cone of $\mathcal{S}$} is a cone of $({\mathbb{R}}_{\geq 0})^{k}$.
@@
More generally, we may, as in Construction 3.3, perform the multivariable construction relatively over the torus $({\rm Sym}_{g\times g}({\mathbb{Z}})/\langle{\mathbb{B}}\rangle)\otimes{\mathbb{C}}^{*}$ by twisting the $\bf M$}-action by some elements $a=(a_{ij})\in{\rm Sym}_{g\times g}({\mathbb{C}}^{*})$. Here, we have quotiented by $\langle{\mathbb{B}}\rangle\otimes{\mathbb{C}}^{*}$, so as to reduce redundant moduli of the general fiber as much as possible. $\clubsuit$
@@
Construction 3.10 (multiparameter case, cone of a fan for ${\mathcal{A}}_{g}$).
@@
is not standard affine, or not even simplicial. For an appropriate choice of fan $\mathcal{S}$}, Construction 3.8 gives a degeneration of abelian varieties over a $k$-dimensional polydisk $X({\mathcal{S}})\to\Delta^{k}\subset Y({\mathbb{R}}_{\geq 0}^{k})$. But in fact, we may descend the construction to an analytic open neighborhood of the torus fixed point in the affine toric variety $Y_{{\mathbb{R}}{\mathbb{B}}}({\mathbb{B}})$, by taking a fan $\mathcal{S}$} supported rather in the vector space ${\bf N}_{\mathbb{R}}\times{\mathbb{R}}{\mathbb{B}}$, and periodic with respect to the same action (12). Here, the subscript ${\mathbb{R}}{\mathbb{B}}$ of $Y_{{\mathbb{R}}{\mathbb{B}}}$ refers to the fact that we take the toric variety of the polyhedral cone ${\mathbb{B}}\subset{\mathbb{R}}{\mathbb{B}}$, sitting inside the vector space ${\mathbb{R}}{\mathbb{B}}$, rather than inside ${\rm Sym}_{g\times g}({\mathbb{R}})$.
Taking the universal twist by $a\in({\rm Sym}_{g\times g}({\mathbb{Z}})/\langle{\mathbb{B}}\rangle)\otimes{\mathbb{C}}^{*}$, we produce a degeneration
$X^{\rm univ}({\mathcal{S}})\to T({\mathbb{B}})\subset Y({\mathbb{B}})$
@@
where now $Y({\mathbb{B}})$ is the toric variety of the polyhedral cone ${\mathbb{B}}\subset{\rm Sym}_{g\times g}({\mathbb{R}})$ and $T({\mathbb{B}})$ is an analytic tubular neighborhood of the deepest toric boundary stratum. Then, following Section 2.5, we may analytically glue $T({\mathbb{B}})$ along the complement $T^{*}({\mathbb{B}})$ of the toric boundary to ${\mathcal{A}}_{g}$ or to an appropriate étale cover $\widetilde{{\mathcal{A}}}_{g}\to{\mathcal{A}}_{g}$. Thus, we produce a degeneration (which we denote by the same symbol)
$X^{\rm univ}({\mathcal{S}})\to\widetilde{{\mathcal{A}}}_{g}\cup_{T^{*}({\mathbb{B}})}T({\mathbb{B}})\subset\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$
which analytically extends the universal family over $\widetilde{{\mathcal{A}}}_{g}$. $\clubsuit$
@@
We have been somewhat casual about the “universal” versions of the Mumford construction, e.g. $X^{\rm univ}({\mathcal{S}})\to\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$. In fact, under the hypothesis that the $B_{i}$ generate a cone in ${\mathcal{P}}_{g}^{+}$, one may unify the set-up of toroidal extensions of ${\mathcal{A}}_{g}$ laid out in Section 2.5, with the toroidal construction of $X({\mathcal{S}})$, by forming a fan in the larger space ${\mathcal{P}}_{g}^{+}\oplus{\mathbb{R}}^{g}\subset{\rm Sym}_{g\times g}({\mathbb{R}})\oplus{\mathbb{R}}^{g}$. Namikawa referred to such fans as “mixed cone decompositions,” see [49, Sec. 3] and [51, Sec. 9] (though in these texts, the focus is one particular Mumford construction, similar to Construction 6.5). We refer to loc. cit. for further details.
@@
Regardless, this finite cover does exist in an affine open neighborhood of the relevant boundary stratum. By a slight abuse of notation, we continue to notate the resulting cover and its toroidal extension by $\widetilde{{\mathcal{A}}}_{g}$ and $\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$, even though $\widetilde{{\mathcal{A}}}_{g}$ is only étale over a Zariski open subset of ${\mathcal{A}}_{g}$.
To summarize Constructions 3.8 and 3.10:
@@
Let $\mathcal{S}$} be a fan in ${\bf N}_{\mathbb{R}}\times{\mathbb{R}}{\mathbb{B}}$ satisfying the properties described in Constructions 3.8 and 3.10. In particular, $\mathcal{S}$} is $\bf M$}-invariant under the action (12), for a collection of symmetric bilinear forms $\{B_{1},\dots,B_{k}\}\subset{\rm Sym}^{2}{\bf M}^{\vee}$ which are the rays generating a polyhedral cone in the space of positive semidefinite bilinear forms on $\bf M$}, with $\sum_{i=1}^{k}B_{i}$ positive-definite.
Then Construction 3.10 produces a flat, proper extension of the universal family over $\widetilde{{\mathcal{A}}}_{g}$, in the category of complex analytic spaces,
$X^{\rm univ}({\mathcal{S}})\to\widetilde{{\mathcal{A}}}_{g}^{+}\subset\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}.$
Here $\widetilde{{\mathcal{A}}}_{g}^{+}$ is the union of $\widetilde{{\mathcal{A}}}_{g}$ with a tubular neighborhood of the deepest toroidal stratum of $\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$.

### v3

@@ 3. The Mumford construction
In Section 2.5, we have shown how to extend ${\mathcal{A}}_{g}$ (or $\widetilde{{\mathcal{A}}}_{g})$ toroidally, to obtain ${\mathcal{A}}_{g}\hookrightarrow{\mathcal{A}}_{g}^{\,{\mathbb{B}}}$ (or $\widetilde{{\mathcal{A}}}_{g}\hookrightarrow\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$), so that the period map $(\Delta^{*})^{k}\to{\mathcal{A}}_{g}$ of any degeneration $f\colon X\to\Delta^{k}$ with monodromy cone ${\mathbb{B}}_{f}={\mathbb{B}}$ admits an extension of the period map over the punctures $\Delta^{k}\to{\mathcal{A}}_{g}^{\,{\mathbb{B}}}$. We will now describe how to extend the universal family of PPAVs over ${\mathcal{A}}_{g}$ so that we may pull back this extension to produce particularly nice birational models of degenerations. It is useful to have examples in mind; many are provided in Section 3.5. All degenerations we consider in this section are maximal, in the sense of Definition 2.13.
@@ 3.1. Mumford construction, fan version
We must take $|u|$ small enough that $B\log|u|-\log|a_{ij}|>0$. Then this construction may be performed relatively, choosing $a$ in some subtorus of ${\rm Sym}_{g\times g}({\mathbb{C}}^{*})$ forming coset representatives for the natural action of $u\in{\mathbb{C}}^{*}$. We get an analytic degeneration
$f^{\rm univ}_{\circ}\colon X^{\rm univ}_{\circ}({\rm Cone}({\mathcal{T}}))\to\Delta^{\rm univ}$
where $\Delta^{\rm univ}\to{\rm Sym}_{g\times g}({\mathbb{C}}^{*})/{\mathbb{C}}^{*}$ is a holomorphic disk bundle over ${\rm Sym}_{g\times g}({\mathbb{C}}^{*})/{\mathbb{C}}^{*}$, such that $f^{\rm univ}_{\circ}$ restricts to a family of PPAVs over the punctured disc bundle $(\Delta^{\ast})^{\rm{univ}}\subset\Delta^{\rm{univ}}$.
In terms of the toroidal extensions of Section 2.5, $\Delta^{\rm univ}$ maps to the toroidal extension ${\mathcal{A}}_{g}^{{\mathfrak{F}}}$ whose fan is (the orbit of) a single ray ${\mathfrak{F}}=\operatorname{GL}_{g}({\mathbb{Z}})\cdot{\mathbb{R}}_{\geq 0}B$. So every $1$-parameter family of PPAVs with monodromy $B$ admits an extension pulled back from $f^{\rm univ}_{\circ}$ along an arc transverse to the boundary divisor $\{0\}\times{\rm Sym}_{g\times g}({\mathbb{C}}^{*})/{\mathbb{C}}^{*}=\Delta^{\rm univ}\setminus(\Delta^{*})^{\rm univ}$. $\clubsuit$
@@
When $\mathcal{T}$} has integral vertices, the total space $X({\rm Cone}({\mathcal{T}}))$ or $X^{\rm univ}_{\circ}({\rm Cone}({\mathcal{T}}))$ of the Mumford degeneration is smooth if and only if $\mathcal{T}$} is a complete triangulation, i.e. $\mathcal{T}$} is a tiling by standard lattice simplices. This condition ensures that ${\rm Cone}({\mathcal{T}})$ is a regular fan—its cones are all standard affine cones. Then $X({\rm Cone}({\mathcal{T}}))\to\Delta$ is a semistable, $K$-trivial degeneration (with smooth total space), see Proposition 3.14 below. These are sometimes called Kulikov models, in analogy to K3 degenerations.
@@
By passing to an intermediate cover ${\mathcal{H}}_{g}\to\widetilde{{\mathcal{A}}}_{g}\to{\mathcal{A}}_{g}$ as in Remark 2.28, we may assume that the Levi quotient $\Gamma\subset\operatorname{GL}_{g}({\mathbb{Z}})$ of the parabolic stabilizer at some Baily–Borel 0-cusp of $\widetilde{{\mathcal{A}}}_{g}$ acts freely on ${\mathcal{P}}_{g}$. Then, when $B$ is primitive, the boundary divisor of $\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{R}}_{\geq 0}B}$ is isomorphic to the discriminant divisor of $f^{\rm univ}_{\circ}$, rather than a further finite quotient of it. In this case, the base of $f^{\rm univ}_{\circ}$ glues onto $\widetilde{{\mathcal{A}}}_{g}$ and thus, we may extend $f^{\rm univ}_{\circ}$ to a family $f^{\rm univ}\colon X^{\rm univ}({\rm Cone}({\mathcal{T}}))\to\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{R}}_{\geq 0}B}$ extending the universal family over $\widetilde{{\mathcal{A}}}_{g}$. A priori, this extension exists only in the category of analytic spaces. But it is always an algebraic space (Prop. 6.12), and for certain choices of tiling $\mathcal{T}$}, we may ensure it is relatively projective, see Section 3.3.
When $B$ is not primitive, the base $\Delta^{\rm univ}$ of $f^{\rm univ}_{\circ}$ rather maps to $\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{R}}_{\geq 0}B}$ by a map ramified over the toroidal boundary divisor, to order the imprimitivity of $B$.
@@
satisfy $\widetilde{\Phi}(\tau)=\widetilde{\Phi}_{\rm nilp}(\tau)=\tau B,$ where $u=e^{2\pi i\tau}$ is the coordinate on $\Delta^{*}$. More generally, the period map on the restriction of $f^{\rm univ}_{\circ}$ to the universal cover of $(\Delta^{*})^{\rm univ}|_{\{a\}}$ for $a\in{\rm Sym}_{g\times g}({\mathbb{C}}^{*})$, is given by $\widetilde{\Phi}(\tau)=\widetilde{\Phi}_{\rm nilp}(\tau)=\tfrac{1}{2\pi i}(\log a_{ij})+\tau B$.
@@
The fan $\mathcal{S}$} must then be $\bf M$}-periodic (with respect to the action (14)) and the projection to ${\mathbb{R}}^{k}$ must induce a morphism of fans to $({\mathbb{R}}_{\geq 0})^{k}$, such that ${\rm Supp}({\mathcal{S}})$ contains ${\mathbb{R}}^{g}\times(1,\dots,1)$. Furthermore, we usually require that the fan morphism ${\mathcal{S}}\to({\mathbb{R}}_{\geq 0})^{k}$ is flat, that is, the image of any cone of $\mathcal{S}$} intersected with ${\bf N}\times{\mathbb{Z}}^{k}$ is a cone of $({\mathbb{R}}_{\geq 0})^{k}$ intersected with ${\mathbb{Z}}^{k}$.
@@
More generally, we may, as in Construction 3.3, perform the multivariable construction relatively over the torus $({\rm Sym}_{g\times g}({\mathbb{Z}})/\langle{\mathbb{B}}\rangle)\otimes{\mathbb{C}}^{*}$ by twisting the $\bf M$}-action by some elements $a=(a_{ij})\in{\rm Sym}_{g\times g}({\mathbb{C}}^{*})$. Here, we have quotiented by $\langle{\mathbb{B}}\rangle\otimes{\mathbb{C}}^{*}$, so as to reduce redundant moduli of the general fiber as much as possible. We denote the resulting fibration by
$f^{\rm univ}_{\circ}\colon X^{\rm univ}_{\circ}({\mathcal{S}})\to(\Delta^{k})^{\rm univ}$
where $(\Delta^{k})^{\rm univ}\to({\rm Sym}_{g\times g}({\mathbb{Z}})/\langle{\mathbb{B}}\rangle)\otimes{\mathbb{C}}^{*}$ is a polydisk bundle with
$\dim(\Delta^{k})^{\rm univ}=g(g+1)/2+k-\dim\langle{\mathbb{B}}\rangle.$
Note that $f^{\rm univ}_{\circ}\colon X^{\rm univ}_{\circ}({\mathcal{S}})\to(\Delta^{k})^{\rm univ}$ is a locally trivial deformation of $f\colon X({\mathcal{S}})\to\Delta^{k}$. Indeed, as analytic germs about the zero section of $(\Delta^{k})^{\rm univ}$, the universal cover of the former is the product of the universal cover of the latter with the torus of twists $({\rm Sym}_{g\times g}({\mathbb{Z}})/\langle{\mathbb{B}}\rangle)\otimes{\mathbb{C}}^{*}$. $\clubsuit$
Remark 3.9.
In terms of fans, the (possibly rational) origin section of $X({\mathcal{S}})\to\Delta^{k}$ is declared to be the image of the subtorus whose cocharacter lattice is $\{0\}\times{\mathbb{Z}}^{k}\subset{\bf N}\times{\mathbb{Z}}^{k}$.
@@
Construction 3.11 (multiparameter case, cone of a fan for ${\mathcal{A}}_{g}$).
@@
is not standard affine, or not even simplicial. Recall from Remark 2.28 that for some étale cover $\widetilde{{\mathcal{A}}}_{g}\to{\mathcal{A}}_{g}$, we have a toroidal extension $\widetilde{{\mathcal{A}}}_{g}\hookrightarrow\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$ whose monodromy cone is $\mathbb{B}$}. For an appropriate choice of fan $\mathcal{S}$}, Construction 3.8 gives a degeneration of abelian varieties over a $k$-dimensional polydisk $X({\mathcal{S}})\to\Delta^{k}\subset Y({\mathbb{R}}_{\geq 0}^{k})$. In fact, we may descend the construction to an analytic open neighborhood of the torus fixed point in the affine toric variety $Y_{{\mathbb{R}}{\mathbb{B}}}({\mathbb{B}})$, when $\mathcal{S}$} is supported rather in the vector space ${\bf N}_{\mathbb{R}}\times{\mathbb{R}}{\mathbb{B}}$, and periodic with respect to the same action (14). Here, the subscript ${\mathbb{R}}{\mathbb{B}}$ of $Y_{{\mathbb{R}}{\mathbb{B}}}$ refers to the fact that we take the toric variety of the polyhedral cone ${\mathbb{B}}\subset{\mathbb{R}}{\mathbb{B}}$, sitting inside the vector space ${\mathbb{R}}{\mathbb{B}}$, rather than inside ${\rm Sym}_{g\times g}({\mathbb{R}})$.
Taking the universal twist, we produce a degeneration
$X^{\rm univ}_{\circ}({\mathcal{S}})\to T({\mathbb{B}})\subset Y({\mathbb{B}})$
@@
where now $Y({\mathbb{B}})$ is the toric variety of the polyhedral cone ${\mathbb{B}}\subset{\rm Sym}_{g\times g}({\mathbb{R}})$ and $T({\mathbb{B}})$ is an analytic tubular neighborhood of the deepest toric boundary stratum. Then, following Section 2.5, we may analytically glue $T({\mathbb{B}})$ along the complement $T^{*}({\mathbb{B}})$ of the toric boundary to $\widetilde{{\mathcal{A}}}_{g}$. Taking the gluing to respect the zero sections, we produce a degeneration
$X^{\rm univ}_{+}({\mathcal{S}})\to\widetilde{{\mathcal{A}}}_{g}\cup_{T^{*}({\mathbb{B}})}T({\mathbb{B}})\eqqcolon\widetilde{{\mathcal{A}}}_{g}^{+}\subset\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$
which analytically extends the universal family $\widetilde{{\mathcal{X}}}_{g}\to\widetilde{{\mathcal{A}}}_{g}$.
One may unify the set-up of toroidal extensions of ${\mathcal{A}}_{g}$ laid out in Section 2.5, with the toroidal construction of $X({\mathcal{S}})$, by viewing $\mathcal{S}$} as a fan ${\mathcal{S}}^{\rm univ}$ in the larger space ${\mathbb{R}}^{g}\oplus{\mathcal{P}}_{g}^{+}\subset{\mathbb{R}}^{g}\oplus{\rm Sym}_{g\times g}({\mathbb{R}})$ admitting a morphism of fans ${\mathcal{S}}^{\rm univ}\to{\mathbb{B}}$ via projection to the second factor. Namikawa was the first to introduce such “mixed cone decompositions”, see [51, Sec. 3] and [53, Sec. 9] (though in these texts, the focus is one particular Mumford construction, similar to Construction 6.5).
The family $X^{\rm univ}_{+}({\mathcal{S}})$ further extends to a proper, flat family
(17) $f^{\rm univ}\colon X^{\rm univ}({\mathcal{S}})\to\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$
surjecting onto the base $\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$. Indeed, this follows from [17, Ch. VI.1]. Defining ${\mathcal{S}}^{\rm univ}$ as above, and taking $T_{\rm max}({\mathbb{B}})\subset Y({\mathbb{B}})$ as the maximal analytic open neighborhood of the toric boundary over which the action of $\bf M$} on the inverse image of $T_{\rm max}({\mathbb{B}})$ in $Y({\mathcal{S}}^{\rm univ})$ is properly discontinuous, the base $T_{\rm max}({\mathbb{B}})$ surjects onto the toroidal extension $\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$. Concretely, $T_{\rm max}({\mathbb{B}})$ is the relative closure of $U_{\mathbb{Z}}\backslash{\mathcal{H}}_{g}\subset{\rm Sym}_{g\times g}({\mathbb{C}}^{*})$, see (8), in the toric variety $Y({\mathbb{B}})$. Then, the family descends from a partial uniformization of toroidal strata fibering over Baily–Borel strata of intermediate dimension, to yield (17).
The precise relation to Faltings–Chai [17, Ch. VI.1] is that the fan ${\mathcal{S}}^{\rm univ}$ is an instance of a $\operatorname{GL}(X)\ltimes X^{s}$-admissible polyhedral cone decomposition” (for $s=1$) in the terminology of loc.cit. up to the following three minor modifications: We do not demand as in [17, Ch. VI, Def.1.3.(iii)] that ${\mathcal{S}}^{\rm univ}$ defines a complete fan (i.e. a compactification of ${\mathcal{X}}_{g}$), when working on an étale cover $\widetilde{{\mathcal{A}}}_{g}\to{\mathcal{A}}_{g}$ we only demand admissibility for a finite index subgroup, and we do not require the existence of a polarization function, see [17, Ch. VI.1, Rem. 1.6.(a)]. The constructions used to prove [17, Ch. VI.1, Thm. 1.13] generalize to this setting, to yield an extension $\widetilde{{\mathcal{X}}}_{g}^{\,{\mathfrak{G}}}\to\widetilde{{\mathcal{A}}}_{g}^{\,{\mathfrak{F}}}$ of the universal family, for any morphism ${\mathfrak{G}}\to{\mathfrak{F}}$ from a mixed cone decomposition (for $\widetilde{{\mathcal{X}}}_{g}$) to a cone decomposition (for $\widetilde{{\mathcal{A}}}_{g}$). The family (17) is a special case.
It is crucial for our applications in [27] to achieve both a smooth total space and flatness over $\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$ (see Proposition 5.3 and Theorem 7.1). As mentioned in [17, Ch. VI, Def. 1.3.(v), Rem. 1.4], achieving both of these properties is a hard combinatorial problem, closely related to the main result of [2]; in general, it is only possible after modifying the base $\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$.
As additional historical notes, the first example of a complete fan for ${\mathcal{X}}_{g}$ (equidimensional but not regular) was provided by Namikawa, see e.g. [51, Sec. 13, Prop. 13.5, Thm. 13.6]. The generalization of Faltings–Chai to the more general setting of mixed Shimura varieties is Pink’s dissertation, see especially [56, Ex. 2.25, Ch. 6, Ch. 10] for discussion relevant to ${\mathcal{X}}_{g}$. $\clubsuit$
[…]
@@
Regardless, this finite cover does exist in an affine open neighborhood of the relevant boundary stratum. By a slight abuse of notation, we continue to notate the resulting cover and its toroidal extension by $\widetilde{{\mathcal{A}}}_{g}$ and $\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}$, even though $\widetilde{{\mathcal{A}}}_{g}$ is only a finite étale over a Zariski open subset of ${\mathcal{A}}_{g}$.
To summarize Constructions 3.8 and 3.11:
@@
Let $\mathcal{S}$} be a fan in ${\bf N}_{\mathbb{R}}\times{\mathbb{R}}{\mathbb{B}}$ satisfying the properties described in Constructions 3.8 and 3.11. In particular, $\mathcal{S}$} is $\bf M$}-invariant under the action (14), for a collection of symmetric bilinear forms $\{B_{1},\dots,B_{k}\}\subset{\rm Sym}^{2}{\bf M}^{\vee}$ which are the rays generating a polyhedral cone in the space of positive semidefinite bilinear forms on $\bf M$}, with $\sum_{i=1}^{k}B_{i}$ positive-definite.
Then Construction 3.11 produces a flat, proper extension of the universal family over $\widetilde{{\mathcal{A}}}_{g}$, in the category of complex analytic spaces,
$X^{\rm univ}({\mathcal{S}})\to\widetilde{{\mathcal{A}}}_{g}^{\,{\mathbb{B}}}.$